# Geometry-Informed Neural Operator Transformer

### v3

@@ 4 Conclusions
In this work, we have presented the Geometry-Informed Neural Operator Transformer (GINOT), a novel framework designed for efficient and accurate solution predictions on arbitrary geometries represented by surface point clouds, without requiring additional geometric features such as signed distance functions (SDFs). By utilizing transformer-based attention mechanisms and a geometry-aware encoding strategy, GINOT effectively captures geometric representations from boundary point clouds and maps them to solution fields at arbitrary query points. Unlike existing neural operator methods constrained by structured meshes, fixed query points, or computationally intensive geometric representations, GINOT offers a flexible and scalable approach for solving complex geometry problems. To enhance its functionality, we have integrated additional encoders to process problem-specific inputs, such as material properties and boundary conditions.
@@ 8 Author contributions statement
Q. Liu conceptualized the study, developed the methodology, implemented the models, prepared the datasets, conducted the analysis, and drafted, reviewed, and edited the manuscript. V. Zhong contributed to running the baseline models (GNO and GI-FNO), manuscript review and editing. H. Meidani contributed to conceptualization, results discussion, and manuscript review and editing. D. Abueidda contributed to results discussion and manuscript review and editing. S. Koric prepared the bracket lug dataset, contributed to conceptualization, results discussion, and reviewed and edited the manuscript. P. Geubelle contributed to the supervision, methodology, results discussion, and manuscript review and editing.

### v4

@@ 4 Conclusions
In this work, we introduced the Geometry-Informed Neural Operator Transformer (GINOT), a novel framework for efficient and accurate solution prediction on arbitrary geometries represented by unordered, non-uniform, and variable-sized surface point clouds, without the need for additional geometric features such as signed distance functions (SDFs). By utilizing transformer-based attention mechanisms and a geometry-aware encoding strategy, GINOT effectively captures geometric representations from boundary point clouds and maps them to solution fields at arbitrary query points. Unlike existing neural operator methods constrained by structured meshes, fixed query points, or computationally intensive geometric representations, GINOT offers a flexible and scalable approach for solving complex geometry problems. To enhance its functionality, we have integrated additional encoders to process problem-specific inputs, such as material properties and boundary conditions.
@@ 8 Author contributions statement
Q. Liu conceptualized the study, developed the methodology, implemented the models, prepared the datasets, conducted the analysis, and drafted, reviewed, and edited the manuscript. W. Zhong contributed to running the baseline models (GNO and GI-FNO), manuscript review and editing. H. Meidani contributed to conceptualization, results discussion, and manuscript review and editing. D. Abueidda contributed to results discussion and manuscript review and editing. S. Koric prepared the bracket lug dataset, contributed to conceptualization, results discussion, and reviewed and edited the manuscript. P. Geubelle contributed to the supervision, methodology, results discussion, and manuscript review and editing.